27 January 2004
LARGE HADRON COLLIDER COMMITTEE
Minutes of the sixty-eighth meeting held on
Wednesday and Thursday, 28-29 January 2004
1. TOTEM Technical Design Report: Karsten
2. Report from the 9th Worskhop on Electronics for the LHC Experiments: Geoff Hall
3. ALICE Status Report: Juergen Schukraft
4. ALICE Trigger/HLT/DAQ/DCS/ECS Technical
Design Report: Orlando Villalobos Baillie, Volker Lindenstruth,
Pierre Vande Vyvre, Lennart Jirden, Sandro Vascotto
*) Part time
The report from the LCG Project Comprehensive Review (LHCC 2003-065 / G-064) was approved without modification.
2. REPORT FROM THE CHIEF SCIENTIFIC OFFICER
The Chief Scientific Officer (CSO) informed the LHCC on various issues concerning CERN in general and the LHC Project in particular.
He underlined the importance of the LHCC as a scientific committee in its role to review the LHC experiments and to serve as an advisory body. This position was reinforced by the Chief Executive Officer.
The LHC Project Leader reported on the status of the LHC Machine. Installation of the machine has been hampered by delays in the cryoline installation in Sector 7-8. Discussions are presently underway with the contractor to find ways to speed up installation and re-scheduling of activities is in progress. The goal remains to close the LHC ring at the end of 2006 so that operation of the machine can commence in April 2007. Further information will be available in March once the cryoline installation in Sector 7-8 is fully underway. Moreover, as can be seen on the LHC Dashboard, construction and delivery of the main dipole magnets is progressing well.
3. REPORT FROM THE TOTEM REFEREE
The Committee heard a report on the refereeís preliminary reactions to the TOTEM Technical Design Report (LHCC 2004-002 / TDR 1). The Technical Design Report describes the method TOTEM proposes to use for the measurement of the total proton-proton cross-section and for the study of elastic scattering and diffractive dissociation at the LHC. The experimental set-up consists of two tracking telescopes T1 and T2 installed on each side of the interaction point to measure the inelastic proton-proton interactions in the forward region, and of two sets of silicon detectors in Roman Pots located symmetrically on each side of the interaction point to measure the elastic cross-section. (An additional Roman Pot station could be equipped later). The T1 telescope consists of five Cathode Strip Chambers while the choice between a Silicon Tracker and a tracker with GEM chambers is being considered for the T2 telescope. Finally, while the measurement of the total cross-section and elastic scattering can be performed using only the TOTEM detectors, the integration of TOTEM into the CMS detector offers the possibility of more detailed studies of diffractive events. The Committee encourages the TOTEM Collaboration to continue developing their experimental design and following further questions and discussions, the referee will make a full report on the TOTEM Technical Design Report at the next meeting of the LHCC.
The referee also reported on the TOTEM request to test the first Roman Pot prototype in the SPS ring in 2004. The referee addressed the outstanding issues from the previous LHCC meeting regarding the implications for the SPS schedule, operation parameters and vacuum, and the detailed plan of measurements TOTEM proposes to carry-out. In view of the resolution of the above open issues and due to the importance of the tests prior to the Roman Pot installation in the LHC, the Committee recommends approval of the tests and requested that the SPS & PS Physics Coordinator to allocate the requested beam time in the SPS schedule.
The first drafts of the 2004 PS East Hall and SPS user schedules have been released. The user requests have been allocated and some beam time is still free. The request for a 2-month neutron irradiation in the PS East Hall behind the DIRAC beam dump depends on the decision to be taken in either April or June on a DIRAC run in 2004. Discussions are ongoing to accommodate the ATLAS EMEC/HEC/FCAL and NA60 requests for beam time on the North Area H6 line of the SPS.
The Coordinator also reported on the request for a 3-day
extension to the machine development period regarding the scrubbing studies in
the SPS. The results of the studies carried-out in 2003 show that the scrubbing
efficiency at cryogenic temperatures are lower than at room temperature.
However, the uncertainties on the measurements on cryogenic surfaces are large
and additional data is needed to better define the scrubbing scenario for the
LHC. Recognising the importance of the scrubbing runs, the LHCC concurred with
the requested 3-day extension to the studies.
5. REPORT ON THE LETTER OF INTENT - `MEASUREMENT OF
PHOTONS AND NEUTRAL PIONS IN THE VERY
FORWARD REGION OF THE LHCí
The LHCC heard a report from Y. Muraki on the Letter of Intent to measure photons and neutral pions in the very forward region of the LHC. The aim of the experiment is to provide knowledge of the energy distribution of particles in the very forward region for the understanding of cosmic ray phenomena. The Collaboration proposes to install a scintillating fiber imaging calorimeter at a forward location about 140 m from an LHC interaction point, the choice of which has not been made. The calorimeter has the objective of identifying photons, and measuring individual photon energies above 100 GeV, the incident particle position and the two-photon invariant mass distribution showing a peak around the neutral pion mass.
The Committee noted that the requested beam conditions are similar to those required for the initial commissioning of the LHC machine in 2007.
The LHCC is reviewing the Letter of Intent and will provide its recommendation in spring 2004.
Good progress was reported on the installation work in the UX15 and USA15 caverns. The installation of the metallic structures, the feet and rails, the trucks for the calorimeter assembly, and the laying of the cable trays in UX15 is underway. Problems with the overhead traveling crane and ventilation contracts have now been resolved. Electrical services and rack installation in USA15 has started and installation of the magnet bus bar system, the magnet power supplies and the safety systems are about to commence.
The Committee noted the delays in the second phase of the BT integration. Slow production of the thermal shields has resulted in a 6-month delay to the start of installation of the first BT coil. The corresponding cold test has now been moved to the period May-June 2004. The LHCC considers the schedule to be critical and will keep monitoring the progress.
Good progress was noted on the calorimeters. Progress on the LAr calorimeters, consisting of the EM Barrel, the EMEC Wheels, the HEC Modules and the FCAL, is good as is that for the corresponding cryostats. Integration of the solenoid magnet in the barrel cryostat is starting. Good progress was also reported on the LAr calorimeter front-end electronics, with the only outstanding concern being the difficulty of the low voltage supplies to meet the specifications. The Tile Calorimeter module instrumentation with optical components is complete and production and assembly of the read-out electronics is now advancing well.
The referees also reported on the status of the Inner Detector. The Production Readiness Review for the Pixel front-end electronics was completed, and series production of the electronics is now underway. Production of the Barrel and End-cap SCT modules is progressing but concerns remain on the Pitch Adapters for the Barrel SCT. The Production Readiness Review for the Inner Detector services was held in December 2003 and the schedule to purchase and install the systems is considered to be tight. Progress was reported on the Web electronic circuits for the TRT End-cap with around 40 good-quality circuits having been delivered. However, construction of the TRT End-cap A is still limited by the shortage of these circuits. Boards for the TRT front-end electronics are being delivered but their rate of delivery still requires careful monitoring. Testing of the ATMEL TRT analogue electronics is ongoing.
Production of all chambers for the Muon Spectrometer is advancing well. CSC production is complete, the MDTs are being equipped with the gas distribution system and Faraday cages at the required rate and most of the TGC chambers have been produced. The Committee, however, did note delays in the RPC production but a reasonable plan is being implemented to recover some of the delay. With the exception of the concern regarding the timely arrival of the Big Wheel structures, all other support structures are being produced on schedule. The LHCC plans to review the status of the Muon Spectrometer front-end electronics at the next meeting of the Committee.
The LHCC reviewed the ATLAS Computing Project. The Collaboration is preparing for the Data Challenge II which will be carried-out during 2004 and will consist of the production of simulated events, their reconstruction and the subsequent test of distributed analysis. Progress was noted on finding solutions to the shortage of manpower for the core software. Moreover, an ATLAS-wide database project is being created. The LHCC urges ATLAS to do this in collaboration with the LCG Project in order to ensure a common solution across all experiments.
The Committee took note of the ATLAS Combined Test Beam runs planned for 2004 at the SPS. A general ATLAS multi-system combined test beam is planned for H8 while a combined EMEC-HEC-FCAL test is planned for the H6 in the North Area.
Finally, the LHCC reviewed the ATLAS milestones and noted
that the number of scheduled and achieved milestones are converging.
7. REPORT FROM THE CMS REFEREES
The LHCC heard a report from the CMS referees, concentrating on progress in the civil engineering, magnet, Tracker, ECAL, and management issues and a preliminary report on the Expression of Interest on diffractive and forward physics.
The LHCC noted the status of the civil engineering and the construction of the magnet. Civil engineering work is continuing at Point 5. Good progress was reported on the magnet with the coupling of module CB-1 to CB-2 being particularly successful. An accident with the lathe at the manufacturer has, however, resulted in a non-recoverable delay of about 6 weeks in the CB+1 module. This will need to be incorporated into the CMS schedule.
The referees also reported on the ECAL. A revised work-plan for the production of the Barrel ECAL crystals was agreed with all partners. The new schedule, nonetheless, has completion of the crystals for the EB- to be about 2.5 months late relative to the CMS Master Schedule V33.2. Also, the plan is based on converting production to 2-in-1 boules which, however, remains very slow to take off. Discussions are ongoing to resolve the above issues and to accelerate production. The first batch of 100 crystals ordered for the End-cap ECAL has been received and is of excellent quality. Good progress was also reported on the discussions with potential additional producers. The ECAL group is awaiting delivery of all final engineering runs for the front-end electronics and in particular a 2-month delay in the ASIC production and packaging was noted. The delay means that the tests in the LHC-type structured beam will now be held in autumn 2004.
The Committee heard a report on the Tracker. The LHCC noted the good progress on the front-end hybrid electronics. Production of the hybrids resumed in November 2003 following qualification of the fix to the cable stiffener and is no longer on the critical path. The LHCC noted as a major concern the delays in production of sensors at ST Microelectronics. Issues related to quality assurance have not been fully resolved between CERN and the firm. The delays threaten completion of the Tracker in time for start of LHC operation in 2007. The LHCC will continue monitoring progress at its next meetings. Module production and testing has started and is advancing satisfactorily.
The referees also reported on the Muon Barrel DTs. Due to a failure of several high voltage boards, installation of the DT and Barrel RPC packages has been suspended pending on-going investigations and a review of the overall high voltage design of the DTs.
The LHCC took note of the CMS Expression of Interest (LHCC-2004-003/G-068) on its intention to enhance the CMS experiment performance in diffractive and forward physics by incorporating the use of the TOTEM detectors and by considering additional detectors to cover the forward region. The referees will make a full report on the Expression of Interest at the next meeting of the LHCC.
Finally, the LHCC reviewed the CMS Master Schedule. The
Collaboration is working on Version v33.4 to accommodate the 6-week additional
delay in the magnet while keeping delays in the DTs in the shadow of the
magnet delay. The ECAL and Tracker remain the most critical projects and the
LHCC will continue monitoring them closely.
8. REPORT FROM THE ALICE REFEREES
The LHCC heard a report from the ALICE referees, concentrating on progress in the sub-systems and a preliminary report on the Trigger, Data Acquisition, High Level Trigger and Control System Technical Design Report (LHCC-2003-062 / TDR 10).
The LHCC took note of the progress in the assembly of the Inner Tracking System (ITS). One full prototype module for the Silicon Pixel Detector (SPD) has been assembled and the bonding efficiency has been improved on the latest ladders. An agreement between ALICE and LHCb on a common sensor production plan is to be presented to the firm VTT for approval. The efficiency of the bonding yields for the Silicon Strip Detector (SSD) is showing marked improvement. Production of front-end hybrid electronics for the Silicon Drift Detector is in progress.
The referees also reported on several of the other sub-systems. Good progress was reported on the TPC, HMPID, Muon Spectrometer, and the PMD. The LHCC noted that the Technical Design Report for the forward detectors is significantly delayed. The TRD supermodule stacking has been delayed by 10 months and no contingency is left in the schedule to have three supermodules complete by autumn 2005. The LHCC requests ALICE to provide a plan to recover the delay.
The LHCC heard a preliminary account on the Trigger, Data Acquisition, High Level Trigger and Control System Technical Design Report. Given that the experiment has been designed primarily to run with heavy-ion beams, the requirements on the lower level trigger selectivity are considered to be modest compared to the other LHC experiments, whereas the trigger complexity is considerable and requires partial or full event reconstruction in the high-level trigger. In addition, a large bandwidth DAQ together with efficient data selection and/or data compression in the High Level Trigger are required to collect sufficient statistics in the short running time available to ALICE. In proton-proton or proton-ion running mode, the requirements on the trigger selectivity are increased and those for the trigger complexity and bandwidth are much reduced. The Control System, consisting of the Detector Control System which has the aim of ensuring the safe and correct operation of the ALICE experiment and the Experiment Control System which is to provide a unified view of the experiment and a central point from where all operations are initiated and controlled, is also described. The LHCC will review the Technical Design Report through the spring and provide its recommendation at the Committeeís session in May 2004.
The next ALICE Comprehensive Review, which will also
incorporate a session on installation of the experiment at Point 2, will be
held at the next LHCC meeting in March.
9. LHCb COMPREHENSIVE REVIEW
The second of the LHCC Comprehensive Reviews of LHCb took place on 26-27 January 2004. The LHCC referees addressed the following areas: Vertex Locator, Inner Tracker and Trigger Tracker, Outer Tracker, RICH Detectors, Calorimeters, Muon System, Trigger, Online, Software and Physics and the topics of Management, Technical Coordination, Integration, Schedules and Costs.
Since the previous Comprehensive Review in January 2003, the LHCb Collaboration has made very significant progress towards the realisation of an experimental set-up ready to record proton-proton collisions at the LHC and the LHCC expects LHCb to have a working detector installed in time for the beginning of LHC operation in April 2007.
In particular, most detector technologies to be used have successfully gone through the R&D phase and construction of the final components is well underway. Installation of the infrastructure and technical services in the experimental area is in progress and issues concerning the interface to the LHC Machine components are being resolved. The LHCC, however, noted as a concern the delay in the assembly and field measurements of the dipole magnet and the slow start-up of production for the Outer Tracker and Muon System.
The LHCC noted, as described in the recommended LHCb Detector Reoptimisation and Trigger Technical Design Reports, the re-optimisation of the LHCb detector has led to a more elegant experiment and with the same physics performance as specified in the Technical Proposal. Moreover, the LHCC took note of the developments in the photodetector for the RICH which have led to the choice of the Hybrid Photo-Diode (HPD).
The principal conclusions and concerns of the LHCC are
summarised below. They will allow the Committee to follow up the outstanding
issues and to monitor future progress of this project in forthcoming sessions
of the LHCC prior to the next LHCb Comprehensive Review one year hence.
30 June - 1 July